OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) Solve the inequality? Show steps? Thank you!

OpenStudy (mathteacher1729):

How far did you get with it before you got stuck?

sam (.sam.):

-2(x-3)>=5-(x+3) -2(x-3)>=5-x-3 -2(x-3)>=2-x -2x+6>=2-x -x+6>=2 -x>=-4 x<=-4*-1 x<=4

OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) First, distribute. But to make it easier you can do this before you distribute \[–2 \times (x + - 3) ≥ 5 (-x + -3)\]

OpenStudy (mathteacher1729):

@CharlesDavid are you interested in seeing how to solve this graphically?